# Covariance test and universal bootstrap by operator norm

### v1

@@ 3.3 Universal bootstrap for covariance test
Theorem 3.5 establishes the uniform Gaussian approximation bound and Type I error bound of the universal bootstrap. Expression (39) ensures that we can uniformly control the test size with an error of at most $n^{-\delta}+B^{-1/2}$, which vanishes as $n\to\infty$ and $B\to\infty$. This result, therefore, guarantees the uniform consistency of the universal bootstrap procedure. To appreciate these results, we compare our universal bootstrap with the widely used multiplier bootstrap that approximates the distribution of $T$ using the distribution of $T^{\text{mb}}=\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}(\bm{X}_{i}\bm{X}_{i}^{T}-\hat{\bm{\Sigma}})\right\|_{\text{op}}$ conditioned on $\bm{X}$, where $\epsilon_{1},\cdots,\epsilon_{n}\in\mathbb{R}$ are random variables with zero mean and unit variance. To ensure consistency of multiplier bootstrap, Han, Xu and Zhou (2018) required a low dimension condition, while Lopes (2022a); Lopes, Erichson and Mahoney (2023); Giessing (2023) imposed a fast eigen-decay condition, as discussed in Section 2. These methods rely on the consistency of sample covariance, i.e. $\|\bm{\hat{\Sigma}}-\bm{\Sigma}\|_{\text{op}}\to 0$, which does not hold even when $\bm{\Sigma}=\bm{I}_{p}$ and $\alpha\geq 1$. This highlights the key advantage of our proposal for covariance testing in a high-dimensional setting where estimation consistency is unattainable, yet test consistency is assured.
3.4 Sharp uniform simultaneous confidence intervals
@@
The approximation of the distribution of the largest singular value has numerous applications, including determining the number of spikes, as discussed in Ding and Yang (2022). Next, we present an additional important application of Theorem 3.5. Define the inner product of two matrices $\bm{A}$ and $\bm{B}$ as $\langle\bm{A},\bm{B}\rangle=\text{tr}(\bm{A}^{T}\bm{B})$, and the Schatten $1-$norm of matrix $\bm{A}$ as $\|\bm{A}\|_{S_{1}}=\sum_{i=1}\sigma_{i}(\bm{A})$. We shall construct sharp uniform simultaneous confidence intervals for $\langle\bm{A},\bm{\Sigma}\rangle$ and $\bm{c}_{1}^{T}\bm{\Sigma}\bm{c}_{2}$ for all $\bm{A}\in\mathbb{R}^{p\times p}$ and $\bm{c}_{1},\bm{c}_{2}\in\mathbb{R}^{p}$. To achieve this, we note that $\hat{q}^{\text{ub},B}_{\bm{\Sigma}}(\alpha)=\hat{q}^{\text{ub},B}_{\text{Sp}(\bm{\Sigma})}(\alpha)$ where the spectral matrix is defined as $\text{Sp}(\bm{\Sigma})=\text{diag}(\lambda_{1}(\bm{\Sigma}),\cdots,\lambda_{p}(\bm{\Sigma}))$. Although the complete matrix $\bm{\Sigma}$ cannot be consistently estimated in our setting, certain estimators $\widehat{\text{Sp}}(\bm{\Sigma})$ of the spectral matrix $\text{Sp}(\bm{\Sigma})$ have been shown to be consistent under a suitable normalized distance. For instance, given the distance $d^{2}_{\text{Sp}}(\widehat{\text{Sp}}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma}))=\text{tr}(\widehat{\text{Sp}}(\bm{\Sigma})-\text{Sp}(\bm{\Sigma}))^{2}/p$, Ledoit and Wolf (2015); Kong and Valiant (2017) proposed estimators $\widehat{\text{Sp}}(\bm{\Sigma})$ satisfying $d_{\text{Sp}}(\widehat{\text{Sp}}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma}))\to 0$ as $n\to\infty$ when $\alpha=1$. Therefore, the estimated threshold $\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma})}(\alpha)$ can replace the unknown $\hat{q}^{\text{ub},B}_{\text{Sp}(\bm{\Sigma})}(\alpha)$.
@@
$\sup_{0\leq\alpha\leq 1}\bigg{|}\mathbb{P}\bigg{(}\forall\bm{A}\in\mathbb{R}^{p\times p},\ \langle\bm{A},\bm{\Sigma}\rangle\in\bigg{[}\langle\bm{A},\bm{\hat{\Sigma}}\rangle-\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma})}(\alpha)$ $\|\bm{A}\|_{S_{1}},$ (40)
$\langle\bm{A},\bm{\hat{\Sigma}}\rangle+\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma})}(\alpha)\ \|\bm{A}\|_{S_{1}}\bigg{]}\bigg{|}\bm{Y}^{1},\cdots,\bm{Y}^{B}\bigg{)}-(1-\alpha)\bigg{|}\lesssim\rho_{n}(\bm{\Sigma}_{0})+$ $d_{\text{Sp}}(\widehat{\text{Sp}}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma}))+B^{-\frac{1}{2}}.$
As a special case, we can also construct sharp uniform simultaneous confidence intervals for $\bm{c}_{1}^{T}\bm{\Sigma}\bm{c}_{2}$
$\sup_{0\leq\alpha\leq 1}\bigg{|}\mathbb{P}\bigg{(}\forall\bm{c}_{1},\bm{c}_{2}\in\mathbb{R}^{p},\ \bm{c}_{1}^{T}\bm{\Sigma}\bm{c}_{2}\in\bigg{[}\bm{c}_{1}^{T}\bm{\hat{\Sigma}}\bm{c}_{2}-\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma})}(\alpha)$ $\|\bm{c}_{1}\|\cdot\|\bm{c}_{2}\|,$ (41)

### v2

@@ 3.2 Consistency of universal bootstrap
Theorem 3.2 establishes the uniform Gaussian approximation bound and Type I error of the universal bootstrap. Expression (13) also ensures that we can uniformly control the test size with an error of at most $n^{-\delta}+B^{-1/2}$, which vanishes as $n\to\infty$ and $B\to\infty$. This result guarantees the uniform consistency of the universal bootstrap procedure.
Remark.
where $T$ is a statistic, $T^{*}$ is its Gaussian counterpart, and $\mathcal{A}$ represents a specified family of sets. For example, Chernozhukov et al. (2013, 2017, 2023) considered $T$ as mean estimators and $\mathcal{A}$ as all rectangular in $\mathbb{R}^{p}$. A more related choice of $\mathcal{A}$ is in Zhai (2018); Xu et al. (2019); Fang & Koike (2024), who also considered mean estimators but take $\mathcal{A}$ to be the sets of Euclidean balls and convex sets in $\mathbb{R}^{p}$. Their results demonstrated that for sets of Euclidean balls $\mathcal{A}$, $\rho(\mathcal{A})$ converge to 00 if and only if $p/n\to 0$, meaning the Gaussian approximation holds when $p=o(n)$. For comparison, we observe that the operator norm ball $\mathbf{B}_{\text{op}}(\bm{v},t)$ for vector $\bm{v}$ in $\mathbb{R}^{p}$ coincides with the Euclidean balls, and our results show that the universality approximation holds when $p/n$ converges to a nonzero constant or even diverges to infinity. For the covariance statistics, Han et al. (2018) took $T$ as the sample covariance matrix and $\mathcal{A}$ as all sets of $s$-sparse operator norm balls (defined in their work). Especially, with $\mathcal{A}$ as all operator norm balls, i.e. $s=p$, they required $p=o(n^{1/9})$, limiting $p$ to be considerably smaller than $n$. Similarly, Lopes (2022b); Lopes et al. (2023) considered $T$ as sample covariance with $\mathcal{A}$ as operator norm balls, but imposed a decay rate $i^{-\beta}$ for the $i$-th largest eigenvalue $\lambda_{i}(\bm{\Sigma})$ of $\bm{\Sigma}$ with $\beta>1$, implying a low intrinsic test dimension. Likewise, Giessing (2023) required the effective rank $r(\bm{\Sigma})=\text{tr}(\bm{\Sigma})/|\bm{\Sigma}|_{\text{op}}$ to satisfy $r(\bm{\Sigma})=o(n^{1/6})$. In contrast, we impose no such assumptions, allowing each eigenvalue of $\bm{\Sigma}$ to be of comparable scale. These comparisons underscore the advantages of our proposal, even in regimes lacking consistency.
3.3 Consistency of generalized universal bootstrap
In this section, we show that the universal bootstrap consistency extends to the general spectral statistic class (4). An important application is the combined statistic $T^{\text{Com}}$ in (5). While Theorem 3.2 provides thresholds for $T$ and $T^{\text{Roy}}$ individually, it cannot be directly applied to $T^{\text{Com}}$ due to the complex dependence between $T$ and $T^{\text{Roy}}$. To address this, we develop the generalized universality theorem. We first introduce
where $\bm{\Sigma}_{2}$ is positive-definite, and $\bm{t}_{k}=(t_{1},\cdots,t_{k})^{T}$ with $t_{i}\geq 0$ for each $i=1,\cdots,k$. Since the operator norm $\|\bm{N}-\bm{\Sigma}_{1}\|_{\text{op}}$ equals the largest singular value $\sigma_{1}(\bm{N}-\bm{\Sigma}_{1})$, we obtain $\mathbf{B}_{\text{op}}(\bm{\Sigma},t)=\mathbf{B}_{1,\text{op}}(\bm{\Sigma},\bm{I}_{p},t)$. Next, we define the following quantity for symmetric matrices $\bm{\Sigma}_{1,m}$ and positive-definite matrices $\bm{\Sigma}_{2,m}$, $m=1,\cdots,M$,
Theorem 3.3 (Generalized universal bootstrap consistency).
Under Assumption 1, given $M$ admissible pairs $(\bm{\Sigma}_{2,m}^{-1/2}\bm{\Sigma}\bm{\Sigma}_{2,m}^{-1/2},\bm{\Sigma}_{2,m}^{-1/2}\bm{\Sigma}_{1,m}\bm{\Sigma}_{2,m}^{-1/2})$ and fixed integer numbers $k_{m}$ for $m=1,\cdots,M$, we have the uniform Gaussian approximation bound for some constant $\delta>0$,
$\rho_{n}(\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M})\lesssim n^{-\delta}\to 0\quad\text{as}\quad n\to\infty.$ (16)
Moreover, we have
Theorem 3.3 extends Theorem 3.2 by demonstrating that universal bootstrap consistency applies to general statistics based on extreme singular values with arbitrary dependencies in the form of (4). This result reinforces our universal bootstrap principle: to approximate the distribution of statistics involving extreme eigenvalues, it suffices to substitute all random variables with their Gaussian counterparts.
4 Application to covariance inference
In this section, we discuss applications of the universal bootstrap in high-dimensional covariance inference. We first apply our universal bootstrapped statistics to the covariance test $H_{0}:\bm{\Sigma}=\bm{\Sigma}_{0}$ with operator norm. As noted in Section 1, several popular statistics utilize the operator norm, including $T^{\text{Roy}}$ in (3). Our Theorem 3.2, 3.3 ensures the size control of universal bootstrap for these statistics. In this subsection, we further perform a power analysis across a family of statistics, with $T$ and $T^{\text{Roy}}$ as specific cases.
We conduct a power analysis under the generalized spike model setting as Jiang & Bai (2021), which allows variability in the bulk eigenvalues and block dependence between the spiked and bulk parts. We assume the following generalized spike structure
where $\bm{D},\bm{V}_{1},\bm{R}_{1}\in\mathbb{R}^{k\times k}$, $\bm{V}_{2},\bm{R}_{2}\in\mathbb{R}^{(p-k)\times(p-k)}$ are diagonal matrices, $k$ is a fixed number, and $\bm{U}$ is orthogonal matrix. We consider the family of statistics $T(\bm{\Sigma}_{1})=\left\|\bm{\Sigma}_{1}^{-\frac{1}{2}}\left(\hat{\bm{\Sigma}}-\bm{\Sigma}_{0}\right)\bm{\Sigma}_{1}^{-\frac{1}{2}}\right\|_{\text{op}}$, and reject $H_{0}$ if $T(\bm{\Sigma}_{1})\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{1}^{-1/2}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-1/2},\bm{\Sigma}_{1}^{-1/2}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-1/2}}(\alpha)$. Notably, $T=T(\bm{I}_{p})$, $T^{\text{Roy}}=T(\bm{\Sigma}_{0})$ are both included in this family. Define $\bm{D}^{\prime}=\bm{D}\bm{R}^{-1}_{1}$, $\bm{V}_{1}^{\prime}=\bm{V}_{1}\bm{R}^{-1}_{1}$, $\bm{V}_{2}^{\prime}=\bm{V}_{2}\bm{R}^{-1}_{2}$. We assume that $\|\bm{D}^{\prime}\|_{\text{op}}$, $\|\bm{V}^{\prime}_{1}\|_{\text{op}}$, $\|\bm{V}^{\prime}_{2}\|_{\text{op}}$ are bounded. Define $E_{+}(\bm{\Sigma}_{1})$, $E_{-}(\bm{\Sigma}_{1})$ and $\tilde{m}(z;\bm{\Sigma}_{1})$ as in Section 3 with $(\bm{\Sigma},\bm{R})=(\bm{V}_{2}^{\prime},-\bm{V}_{2}^{\prime})$ in $\bm{M}_{n}$. For simplicity, we only consider $\gamma=1$.
Theorem 4.1.
Under Assumptions 1, define $\bm{D}^{\prime}=\text{diag}\left((d_{i}^{\prime})_{i=1}^{k}\right)$, $\bm{V}_{1}^{\prime}=\text{diag}\left((v_{i}^{\prime})_{i=1}^{k}\right)$ and
$\kappa=-\frac{\phi^{\frac{1}{2}}}{\tilde{m}(\phi^{-\frac{1}{2}}E_{+}(\bm{\Sigma}_{1});\bm{\Sigma}_{1})}-\frac{\phi^{\frac{1}{2}}v_{i}^{\prime}}{E_{+}(\bm{\Sigma}_{1})\tilde{m}(\phi^{-\frac{1}{2}}E_{+}(\bm{\Sigma}_{1});\bm{\Sigma}_{1})}.$ (25)
(i) If $d_{i}^{\prime}>\kappa$ for some $1\leq i\leq k$, then the power goes to $1$, i.e.
$\mathbb{P}\left(T(\bm{\Sigma}_{1})\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{1}^{-1/2}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-1/2},\bm{\Sigma}_{1}^{-1/2}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-1/2}}(\alpha)\right)\to 1,\ \text{as}\ n\to\infty,\ B\to\infty.$
(ii) If $d_{i}^{\prime}<\kappa$ for all $1\leq i\leq k$ and $T(\bm{\Sigma}_{1})=\lambda_{1}\left(\bm{\Sigma}_{1}^{-1/2}\left(\hat{\bm{\Sigma}}-\bm{\Sigma}_{0}\right)\bm{\Sigma}_{1}^{-1/2}\right)$, then
$\mathbb{P}\left(T(\bm{\Sigma}_{1})\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{1}^{-1/2}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-1/2},\bm{\Sigma}_{1}^{-1/2}\bm{\Sigma}_{0}\bm{\Sigma}_{1}^{-1/2}}(\alpha)\right)\to\alpha,\ \text{as}\ n\to\infty,\ B\to\infty,$
i.e. there is no power under this setting.
Theorem 4.1 provides the power analysis of the test based on the statistic $T(\bm{\Sigma}_{1})$. Specifically, it identifies the sharp phase transition point $\kappa$ in (25), which closely resembles the well-known Baik-Ben Arous-Péché (BBP) phase transition for the largest eigenvalues of the sample covariance matrix, see Baik et al. (2005); Baik & Silverstein (2006); Paul (2007). When the spike eigenvalues exceed this phase transition point, they lie outside the bulk eigenvalue support, yielding full test power. Conversely, if the spike eigenvalues $d^{\prime}_{i}$ fall below this point, distinguishing them from the null case using $T(\bm{\Sigma}_{1})$ becomes infeasible.
Corollary 2.
Under assumptions of Theorem 4.1, define $\bm{D}=\text{diag}\left((d_{i})_{i=1}^{k}\right)$, $\bm{V}_{1}=\text{diag}\left((v_{i})_{i=1}^{k}\right)$.
(i) If $d_{i}-v_{i}>-\frac{\phi^{1/2}}{\tilde{m}(\phi^{-1/2}E_{+})}-v_{i}\left(1+\frac{\phi^{1/2}}{E_{+}\tilde{m}(\phi^{-1/2}E_{+})}\right)$ for some $1\leq i\leq k$, then the power of $T$ goes to $1$. Here $\phi^{-1/2}E_{+}\tilde{m}(\phi^{-1/2}E_{+})<-1$.
(ii) If $d_{i}-v_{i}>\phi^{1/2}v_{i}$ for some $1\leq i\leq k$, then the power of $T^{\text{Roy}}$ goes to $1$.
We compare the phase transition points (i) and (ii) for $T$ and $T^{\text{Roy}}$, respectively. These points reveal distinct behaviors: as $\phi^{-1/2}E_{+}\tilde{m}(\phi^{-1/2}E_{+})<-1$, the phase transition point of (i) decreases with $v_{i}$, whereas the phase transition point of (ii) increases with $v_{i}$. This indicates that when spikes affect the larger eigenvalues, i.e. for large $v_{i}$, the statistics $T$ achieves high power more effectively than the normalized $T^{\text{Roy}}$ with a lower phase transition point. Conversely, the normalized statistic $T^{\text{Roy}}$ performs better for smaller $v_{i}$.
@@ 4 Application to covariance inference
As another important application of the universal bootstrap, we show how to construct sharp uniform simultaneous confidence intervals for $\langle\bm{A},\bm{\Sigma}\rangle$ for all $\bm{A}\in\mathbb{R}^{p\times p}$, where the inner product of two matrices $\bm{A}$ and $\bm{B}$ is defined as $\langle\bm{A},\bm{B}\rangle=\text{tr}(\bm{A}^{T}\bm{B})$. Notably, setting $\bm{A}=\bm{c}_{1}\bm{c}_{1}^{T}$ and $\bm{A}=\bm{c}_{1}\bm{c}_{2}^{T}$ gives the variance $\text{Var}(\bm{c}_{1}^{T}\bm{X}_{i})$ and covariance $\text{Cov}(\bm{c}_{1}^{T}\bm{X}_{i},\bm{c}_{2}^{T}\bm{X}_{i})$ of all linear combinations $\bm{c}_{1}^{T}\bm{X}_{i}$ and $\bm{c}_{2}^{T}\bm{X}_{i}$ of $\bm{X}_{i}$, which has significant applications in optimal portfolio analysis (Bai & Silverstein, 2010; Hult et al., 2012). The main challenge comes from that the full matrix $\bm{\Sigma}$ and consequently $\hat{q}^{\text{ub},B}_{\bm{\Sigma},\bm{\Sigma}}(\alpha)$ cannot be consistently estimated in our setting, so the universal bootstrap cannot be directly applied. To address this, we note that $\hat{q}^{\text{ub},B}_{\bm{\Sigma},\bm{\Sigma}}(\alpha)=\hat{q}^{\text{ub},B}_{\text{Sp}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma})}(\alpha)$ where the spectral matrix is defined as $\text{Sp}(\bm{\Sigma})=\text{diag}(\lambda_{1}(\bm{\Sigma}),\cdots,\lambda_{p}(\bm{\Sigma}))$, and certain estimators $\widehat{\text{Sp}}(\bm{\Sigma})$ for the spectral matrix $\text{Sp}(\bm{\Sigma})$ are consistent under a suitable normalized distance. For instance, given the distance $d^{2}_{\text{Sp}}(\widehat{\text{Sp}}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma}))=\text{tr}(\widehat{\text{Sp}}(\bm{\Sigma})-\text{Sp}(\bm{\Sigma}))^{2}/p$, Ledoit & Wolf (2015); Kong & Valiant (2017) proposed estimators $\widehat{\text{Sp}}(\bm{\Sigma})$ satisfying $d_{\text{Sp}}(\widehat{\text{Sp}}(\bm{\Sigma}),\text{Sp}(\bm{\Sigma}))\to 0$ as $n\to\infty$ when $\gamma=1$.
@@
$\sup_{0\leq\alpha\leq 1}\bigg{|}\mathbb{P}\bigg{(}\forall\bm{A}\in\mathbb{R}^{p\times p},\ \langle\bm{A},\bm{\Sigma}\rangle\in\bigg{[}\langle\bm{A},\bm{\hat{\Sigma}}\rangle-\hat{q}^{\text{ub},B}_{\widehat{\text{Sp}}(\bm{\Sigma}),\widehat{\text{Sp}}(\bm{\Sigma})}(\alpha)$ $\|\bm{A}\|_{S_{1}},\langle\bm{A},\bm{\hat{\Sigma}}\rangle+$ (26)